OpenStudy (anonymous):

Using complete sentences, write an indirect proof proving that if x=30, then 3/4x+5≠20 (3/4 being three over four) I am awful at proofs and worse at indirect proofs. Can someone help me out? And hopefully explain it?

OpenStudy (ikram002p):

first step is to define the given hypothesis to make yo understand the question if x=30, then \(\frac{3}{4}x+5≠20\) so you have Q:- x=30 P:- \(\frac{3}{4}x+5≠20\) and wanna prove if Q then P (Q \(\rightarrow\)P)

OpenStudy (ikram002p):

fine so far ? now there is two type of proving:- type1 : direct proof type 2 : indirect prove like (contrapositive and contradiction )

OpenStudy (anonymous):

Okay I think so. What I understood for the first step was: To prove if ' x equals 30 ' then ' three fourth x plus five does not equal 20 ' ? x=30 being Q and 3/4x+5≠20 being P? And okay so 2 types. Direct is proving it to be true correct? or no?

OpenStudy (ikram002p):

ok , seems u got the point now what proof you wanna me to continue with you contradiction or contrapositive ?

OpenStudy (anonymous):

Can you please do? I don't know what I'm trying to do here. What am I explaining and how should I explain it? I kind of understand what is being showed, I just don't know what to do with what Is being showed.

OpenStudy (ikram002p):

ok ok ^^ since u got what i said above lets prove using contrapositive , could you write the truth table of \(Q \rightarrow P\) ?

OpenStudy (anonymous):

Erm, somewhat? Again I am a lost cause when it comes to proofs. [Sorry about that. Thats why I need help]

OpenStudy (ikram002p):

familour with this ?